OpenStudy (anonymous):

What are the foci of the hyperbola given by the equation (x+10)^2/25 - (y+1)^2/9= 1?

OpenStudy (anonymous):

\[\frac{ (x+10)^2 }{ 25 } - \frac{ (y+1)^2 }{ 9 } = 1\]

OpenStudy (anonymous):

^^^What are the foci of the hyperbola given by the equation??^^^

OpenStudy (anonymous):

first find the center the center of \[\frac{ (x-h)^2 }{ a^2 } - \frac{ (y-k)^2 }{ b^2 } = 1\] is \((h,k)\)

OpenStudy (anonymous):

then the foci are found to be \(\sqrt{a^2+b^2}\) units to the left and right of the center

OpenStudy (anonymous):

in your case \(\sqrt{a^2+b^2}=\sqrt{25+9}=\sqrt{34}\)

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

Sorta, these are the answer choices, (-1 ±√34 , -10) (-1, -10 ± √34 ) (-10 ± √34 , -1) (-10, -1 ± √34)

OpenStudy (anonymous):

as usual, it is C, because it is always C

OpenStudy (anonymous):

center is \((-10,-1)\) go \(\sqrt{34}\) units left and right, you get the answer in C

OpenStudy (anonymous):

Thank you! But can you answer another question of mines?

OpenStudy (anonymous):

What is the eccentricity of the hyperbola given by the equation (y^2/100) - (x^2/36)= 1?